OpenStudy (anonymous):

perform the indicated operations and , if possible simplify. Please show all of work. -7/12 - 1/3

OpenStudy (gw2011):

You need to find the common denominator which in this case is 12. -7/12 = -7/12 -1/3 = -4/12 Add both to get -11/12

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

can u help me with another prblem!

OpenStudy (gw2011):

I'll try

OpenStudy (anonymous):

ok Perform the indicated operation and, if possible, simplify. Please show all of your work. -15/28 * (-4/25)*(3)

OpenStudy (gw2011):

(15/28)(-4/25)(3) To make this problem easier to work with you can cancel some numbers. 15 and 25 can be simplified by dividing 15 by 5 to get 3 and 25 divided by 5 gives 5. -4 and 28 can be simplified by dividing by 4: -4 divided by 4 gives -1 and 28 divided by 4 gives 7. Now the problem looks like this: (3/7)(-1/5)(3) = -9/35 which is the answer

OpenStudy (anonymous):

thanks again!

OpenStudy (gw2011):

You're welcome